# Quality Over Quantity? LLM-Based Curation for a Data-Efficient Audio–Video Foundation Model

### v3

@@ I Introduction and Motivation
Humans naturally process audiovisual information without any need for textual mediation. For instance, when watching a video, we instinctively merge visual cues with corresponding sounds to create a cohesive understanding of the scene. However, most current multimodal AI systems, like CLIP [1] and CLAP [2], and majority of other models [3, 4, 5, 6, 7, 8, 9], depend on textual captions to connect visual and auditory features. This reliance on text-based alignment is at odds with how humans integrate sensory information, where no explicit textual representation is required.
Figure 1: Overview of the proposed audiovisual alignment approach. (a) Our method’s data curation stage uses multimodal reasoning to retain only highly aligned data. It uses LLaVA-Next-Video [10, 11] for video reasoning, LTU-AS [12] for audio processing, and Mistral [13] for alignment scoring. (b) AVVA employs Whisper (audio), DINOv2 (video backbone), without the need for textual mediation during training.
Replicating this human-like processing in AI is challenging [14, 15, 16]. Existing multimodal models primarily handle individual modalities separately, later merging them based on text-based associations [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. This approach, evident in models like Wav2CLIP [5], AudioCLIP [6] and ImageBind [7], misses the opportunity to exploit the natural synchronization between audio and visual data. While efforts like AV-HuBERT [27] and DenseAV [28] aim to capture linguistic information along with the location of sounds from raw audiovisual pairs, they still rely on speech-image pairs for training, which may restrict their generalization.
To address this gap, we introduce AVVA: Audio-Video Vector Alignment, a framework designed to directly align AV information without any text dependency. The proposed model leverages Whisper [29] for audio processing and DINOv2 [30] for visual understanding. This makes AVVA particularly effective in applications requiring concurrent, text-free audiovisual comprehension, such as video analysis and human-computer interaction. An important data curation stage takes place first, which itself relies on a text-, audio- and video-LLMs.
@@ II-A Multimodal Reasoning Engine (MRE)
The potential of most AI techniques for LLMs and multimodal learning often hinges on the diversity and quality of the data they interact with [31, 32]. In this work, we curate the training data via the introduction of a Multimodal Reasoning Engine (MRE), which is a set of prompts for obtaining detailed reasoning of audio, video, and finally to score the level of alignment between audio and video from their textual descriptions, given a set of five metrics.
We used multiple audiovisual datasets that cover diverse scenes from both egocentric and exocentric perspectives, and various forms of audio, including natural, music, ambient, and speech, and other complex scenarios. The datasets are: Epic-Kitchens (1.37 hrs) [33], HowTo100M (7.77 hrs) [34], Music-MIT (2.14 hrs) [35], VALOR (train/test 94.57/13.58 hrs) [36], VGGSound (train/test 30.23/2.82 hrs) [37], AVE (10.00 hrs) [38], AudioSet (54.83 hrs) [39], AudioCaps (train/test 32.64/1.00 hrs) [40], HD-VILA-100M (51.45 hrs) [41]. All input videos were segmented into 3-sec clips; for longer videos, up to 20 random clips were kept. To achieve joint audio-speech reasoning, each segment was processed using LLaVa-NeXT-Video with LLaMA 3 [10, 11] for video reasoning and LTU-AS [12] with LLaMA 2 for audio reasoning. We used Mistral 7B Instruct v0.3 for prompting and measuring alignment.
@@ II-B Model Architecture and Language-free Training
The AVVA model employs a bidirectional cross-modal attention mechanism to integrate audio and video modalities using dual encoders—Whisper for audio and DINOv2 for video. We selected DINOv2 [30] over models like CLIP [1] due to its ability to capture local visual features, which are crucial for producing high-quality global representations through feature pooling [28]. For the audio encoder, we utilize Whisper, concatenating 32 layers while discarding the first layer, as applied in [12]. The architecture is illustrated in Fig. 3.
The bidirectional attention mechanism, implemented with 8 attention heads and a 768-dimensional hidden state, ensures a robust flow of information between the audio and video streams, treating both modalities as equally important. This design makes AVVA particularly effective for complex multimodal tasks requiring detailed audiovisual understanding, such as synchronized multimedia content generation [42], event detection [43], and other tasks requiring detailed audiovisual analysis [44]. By aligning the features through learnable aligner layers - implemented as MLPs with dimensions (input_dim, 1024, 768) and layer normalization, $ReLU$, and dropout 0.2 between layers - and pooling the outputs, the model generates compact embeddings suitable for contrastive learning.
We use the InfoNCE loss function [1] with a temperature setting of 0.07 to help the model learn correlations between audio and video features. For optimization, we adopt the AdamW optimizer with a learning rate and weight decay set to $10^{-4}$. To maintain computational efficiency and prevent catastrophic forgetting, both the DINOv2 and Whisper encoders are frozen [45] during training, focusing on training only the alignment and cross-modal layers.
@@ III-A Cross-modal Retrieval
In this experiment, we assess the ability of AVVA to retrieve audio from video input and vice versa, across three datasets: AudioCaps [40], VALOR [36], and VGG-Sound [37]. Each test is performed on 3-second video segments containing embedded audio, and we compare our model against Wav2CLIP [5], DenseAV [28], Random weights, and ImageBind [7]. Each retrieval test is evaluated K=100 times on N=100 random audio/video files per iteration. No duplication occurs within each set of 100 samples per run. Results are reported as statistical averages.
@@ III-C Temporal Alignment
To investigate the audio-video temporal alignment, we conducted simulations where audio segments were systematically shifted relative to their corresponding video segments across a defined range of shifts from -3.0 to 3.0 sec, in increments of 0.4 sec. For each shift, cosine similarity between audio and video embeddings was computed to assess the alignment quality. The multimodal audiovisual embeddings were extracted using our pre-trained model.
Figure 5 shows cosine similarity scores for the video and audio embeddings of 200 samples across audio shifts. Mean similarity values at each shift are represented by markers, with a smoothed trend line fitted using a Savitzky-Golay filter to highlight the underlying pattern while preserving key variations. The analysis reveals a clear peak in similarity at around 0 sec shift between audio and video, providing evidence of meaningful audio-video learning.
Figure 5: Cosine similarity between AVVA embeddings (video versus shifted audio) as a function of audio shift. The data points show mean similarity scores at each shift level.

### v4

@@ I Introduction and Motivation
Humans naturally process audiovisual information without any need for textual mediation. For instance, when watching a video, we instinctively merge visual cues with corresponding sounds to create a cohesive understanding of the scene. However, most current multimodal AI systems, like CLIP [radford2021learning] and CLAP [elizalde2023clap], and majority of other models [li2023blip, chen2020uniter, wu2022wav2clip, guzhov2022audioclip, girdhar2023imagebind, vosoughi2024learning, chen2023dialogmcf], depend on textual captions to connect visual and auditory features. This reliance on text-based alignment is at odds with how humans integrate sensory information, where no explicit textual representation is required.
Figure 1: Overview of the proposed audiovisual alignment approach. (a) Our method’s data curation stage uses multimodal reasoning to retain only highly aligned data. It uses LLaVA-Next-Video [liu2024llavanext, zhang2024llavanextvideo] for video reasoning, LTU-AS [gong2023joint-ltuas] for audio processing, and Mistral [mistral2023] for alignment scoring. (b) AVVA employs Whisper (audio), DINOv2 (video backbone), without the need for textual mediation during training.
Replicating this human-like processing in AI is challenging [harwath2016unsupervised, tian2020unified, harwath2018jointly]. Existing multimodal models primarily handle individual modalities separately, later merging them based on text-based associations [tang2024video, bi2024eagle, nguyen2024oscar, su2024from, tangsalmonn2024iclr-salmon, kong2024audio, deshmukh2023pengi, gong2024listen, huang2024vtimellm, zhu2024languagebind]. This approach, evident in models like Wav2CLIP [wu2022wav2clip], AudioCLIP [guzhov2022audioclip] and ImageBind [girdhar2023imagebind], misses the opportunity to exploit the natural synchronization between audio and visual data. While efforts like AV-HuBERT [shi2022learning-avhubert] and DenseAV [hamilton2024separating] aim to capture linguistic information along with the location of sounds from raw audiovisual pairs, they still rely on speech-image pairs for training, which may restrict their generalization.
To address this gap, we introduce AVVA: Audio-Video Vector Alignment, a framework designed to directly align AV information without any text dependency. The proposed model leverages Whisper [radford2023whisper] for audio processing and DINOv2 [oquab2023dinov2] for visual understanding. This makes AVVA particularly effective in applications requiring concurrent, text-free audiovisual comprehension, such as video analysis and human-computer interaction. An important data curation stage takes place first, which itself relies on a text-, audio- and video-LLMs.
@@ II-A Multimodal Reasoning Engine (MRE)
The potential of most AI techniques for LLMs and multimodal learning often hinges on the diversity and quality of the data they interact with [sorscher2022beyond, gunasekar2023Phitextbooks]. In this work, we curate the training data via the introduction of a Multimodal Reasoning Engine (MRE), which is a set of prompts for obtaining detailed reasoning of audio, video, and finally to score the level of alignment between audio and video from their textual descriptions, given a set of five metrics.
We used multiple audiovisual datasets that cover diverse scenes from both egocentric and exocentric perspectives, and various forms of audio, including natural, music, ambient, and speech, and other complex scenarios. The datasets are: Epic-Kitchens (1.37 hrs) [damen2020epic], HowTo100M (7.77 hrs) [miech2019howto100m], Music-MIT (2.14 hrs) [zhao2019sound], VALOR (train/test 94.57/13.58 hrs) [chen2023valor], VGGSound (train/test 30.23/2.82 hrs) [chen2020vggsound], AVE (10.00 hrs) [tian2018ave], AudioSet (54.83 hrs) [audioset], AudioCaps (train/test 32.64/1.00 hrs) [kim2019audiocaps], HD-VILA-100M (51.45 hrs) [xue2022hdvila]. All input videos were segmented into 3-sec clips; for longer videos, up to 20 random clips were kept. To achieve joint audio-speech reasoning, each segment was processed using LLaVa-NeXT-Video with LLaMA 3 [liu2024llavanext, zhang2024llavanextvideo] for video reasoning and LTU-AS [gong2023joint-ltuas] with LLaMA 2 for audio reasoning. We used Mistral 7B Instruct v0.3 for prompting and measuring alignment.
@@ II-B Model Architecture and Language-free Training
The AVVA model employs a bidirectional cross-modal attention mechanism to integrate audio and video modalities using dual encoders—Whisper for audio and DINOv2 for video. We selected DINOv2 [oquab2023dinov2] over models like CLIP [radford2021learning] due to its ability to capture local visual features, which are crucial for producing high-quality global representations through feature pooling [hamilton2024separating]. For the audio encoder, we utilize Whisper, concatenating 32 layers while discarding the first layer, as applied in [gong2023joint-ltuas]. The architecture is illustrated in Fig. 3.
The bidirectional attention mechanism, implemented with 8 attention heads and a 768-dimensional hidden state, ensures a robust flow of information between the audio and video streams, treating both modalities as equally important. This design makes AVVA particularly effective for complex multimodal tasks requiring detailed audiovisual understanding, such as synchronized multimedia content generation [xing2024seeing], event detection [berghi2024fusion], and other tasks requiring detailed audiovisual analysis [su2023separating]. By aligning the features through learnable aligner layers - implemented as MLPs with dimensions (input_dim, 1024, 768) and layer normalization, $ReLU$, and dropout 0.2 between layers - and pooling the outputs, the model generates compact embeddings suitable for contrastive learning.
We use the InfoNCE loss function [radford2021learning] with a temperature setting of 0.07 to help the model learn correlations between audio and video features. For optimization, we adopt the AdamW optimizer with a learning rate and weight decay set to $10^{-4}$. To maintain computational efficiency and prevent catastrophic forgetting, both the DINOv2 and Whisper encoders are frozen [wang2023makes] during training, focusing on training only the alignment and cross-modal layers.
@@ III-A Cross-modal Retrieval
In this experiment, we assess the ability of AVVA to retrieve audio from video input and vice versa, across three datasets: AudioCaps [kim2019audiocaps], VALOR [chen2023valor], and VGG-Sound [chen2020vggsound]. Each test is performed on 3-second video segments containing embedded audio, and we compare our model against Wav2CLIP [wu2022wav2clip], DenseAV [hamilton2024separating], Random weights, and ImageBind [girdhar2023imagebind]. Each retrieval test is evaluated K=100 times on N=100 random audio/video files per iteration. No duplication occurs within each set of 100 samples per run. Results are reported as statistical averages.
@@ III-C Temporal Alignment
To investigate the audio-video temporal alignment, we conducted simulations where audio segments were systematically shifted relative to their corresponding video segments across 3-sec increments. For each shift, cosine similarity between audio and video embeddings was computed to assess the alignment quality. The multimodal audiovisual embeddings were extracted using our pre-trained model. Figure 5 shows average cosine similarity scores for the video and audio embeddings of 200 samples across both audio and video shifts. The analysis reveals a clear peak in similarity at 0 sec shift between audio and video, providing evidence of meaningful audio-video learning.
Figure 5: Cosine similarity between AVVA embeddings as a function of audio shifts and video shifts. The data points show mean similarity scores at each shift level, comparing audio embeddings against shifted audio (panel 1) and shifted video (panel 4); and vice versa, similarity of video embeddings against shifted video (panel 2) and shifted audio (panel 3).